# Benoit Decerf [^1] World Bank    Guillaume Haeringer [^2] Baruch College    Martin Van der Linden [^3] Independent scholar

### 07 Jan 2024 v1

@@ 3.2 Incontestable assignments
Our first question is to characterize incontestable assignments under ex-post canonical information. Clearly, incontestable assignments are necessarily individually rational and non-wastefulness. Indeed, given an ex-post canonical information $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q,\mu(i),(|\mu(s)|)_{s\in S}\}$, for any compatible problem $\Gamma=(I,S,\widehat{\succ},r,q)$ and assignment $\widehat{\mu}$ we have ^$\widehat{\succ}_{i}=\succ_{i}$, $\widehat{\mu}(i)=\mu(i)$, and $|\widehat{\mu}(s)|=|\mu(s)|$ for each $s\in S$. So, $s\succ_{i}\mu(i)$ implies $s\succ_{i}\widehat{\mu}(i)$. Also, $s\succ_{i}\mu(i)$ and $|\mu(s)|<q_{s}$ imply $s~{}\widehat{\succ}_{i}~{}\widehat{\mu}(i)$ and $|\widehat{\mu}(s)|<q_{s}$.
@@ 4.1 Incontestable outcomes
Consider the following mechanism $\varphi$. For any profile of preferences $\succ$, $\varphi$ first assigns all students who have a top-priority set (to a school in their top-priority set). Then, $\varphi$ consists of a serial dictatorship with the student $i_{1}$ choosing first if $i_{1}$ does not have a top-priority set. With the preferences we have specified for $i_{1}$, there is no preference profile for students $i_{2}$, $i_{3}$, and $i_{4}$ such that $\varphi$ assigns $i_{3}$ to $s_{3}$. If $i_{2}$ has a top-priority set (which happens whenever she finds at least one school acceptable), $i_{3}$ ends up being assigned to either $s_{1}$ or $s_{2}$. Otherwise, $i_{3}$ is assigned to $s_{1}$.
@@
$\varphi_{j}(\Gamma)=\varphi_{j}(\Gamma^{\prime}),\quad\text{ for all }j\neq i\,,$ (1)
@@
Theorem 2 also establishes an equivalence between the interim perspective and the ex-post perspective (given by the lef-hand side and right-hand side of Eq. (2), respectively). For a student without any top-priority set, any individually rational outcome is a priori (interim) possible, and for any such outcome there exists a preference profile of the other students that rationalizes it (ex-post).
@@ 4.2 Stability and efficiency
Corollary 5 implies that all the mechanisms that Pareto improve the Student-Optimal Stable mechanism (e.g., Kesten,’s (2010) efficiency adjusted mechanism) are incontestable. We provide more details on this later in this section.
@@ 4.3 Contestable mechanisms
Because respect for top-priority sets has bite only for the students who have a top-priority set, the set of incontestable assignments is likely to be significantly larger than the set of stable assignments. But incontestability is not an ‘anything goes’ concept: several well-known assignment mechanisms are contestable. We provide three examples.
@@ 5 Incentives
Theorem 2 has several important implications regarding students’ strategic incentives when participating in an incontestable (and top-top consistent) mechanism.
Incontestability does not guarantee strategyproofness. This is immediate given that there are incontestable mechanisms that are not strategyproof (e.g., EADAM). In spite of this, incontestable mechanisms provide students with some straightforward incentives. We review in this section three types of incentive for assignment mechanisms: maxmin strategies, safe strategies, and manipulability comparisons.
@@ 5.1 Safe strategies
For a student $i$ and a profile of priority rankings and capacities $(r,q)$, we say that a strategy $\succ_{i}$ is safe under mechanism $\varphi$ if student $i$ is guaranteed to be assigned to a school for any preference profile of the other students. Formally, $\succ_{i}$ is safe for student $i$ if
@@ 5.2 Maxmin strategies
Following Reny, (2022), given a mechanism $\varphi$ and an interim canonical information set $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q\}$, let $w_{i}(\succ_{i}^{\prime}|\succ_{i},\varphi,\mathcal{H}_{i})$ be the worst school according to $\succ_{i}$ when student $i$ submits the preference ordering $\succ_{i}^{\prime}$ when participating to the mechanism. Formally, $w_{i}(\cdot)$ is defined as follows,
@@
Theorem 5 of (Reny,, 2022) shows that under the priority efficient mechanism (or EADAM) truth-telling is a maxmin optimal strategy for every student. We can easily generalize Reny’s theorem to any incontestable and top-top consistent mechanism.
@@ 5.3 Dominant strategies
Comparisons in terms of dominant strategies inform about the extent to which mechanisms are manipulable. Over the past decades, some school districts reformed their assignment procedures in the objective to reduce their vulnerability to manipulation (Pathak and Sönmez,, 2013). A mechanism that is less vulnerable to manipulations ‘levels the playing field’ because it should a priori penalizes less students who do not strategize well. Recall that strategyproof mechanisms do not necessarily eliminate the incentive to misrepresent one’s preferences because their constrained version often used in practice is no longer strategyproof —see Haeringer and Klijn, (2009).
Lemma 11.
@@
(i)  The characterization of dominant strategies provided in part (iii) of Lemma 11 does not depend on the specific mechanism. So a student has a dominant strategy in $\varphi^{k}$ if, and only if she has a a dominant strategy in $\psi^{k}$.
(ii)  We first show that $\varphi^{k}$ is less manipulable than $\varphi^{k-1}$. To this end, suppose that student $i$ has a dominant strategy with $\varphi^{k-1}$. By Lemma 11, either $i$ has no more than $k-1$ acceptable schools or $i$’s $k-1$ most preferred schools constitute a top-priority set. These two properties still hold under $\varphi^{k}$, so by Lemma 11 $i$ has a dominant strategy with $\varphi^{k}$.
Consider now the problem $(I,S,\succ,r,q)$, where $q_{s}=1$ for each $s\in S$, and for any two schools $s,s^{\prime}\in S$ and any student $i\in I$, $r_{s}(i)=r_{s^{\prime}}(i)$. Let $i$ be the student for whom exactly $k-1$ other students have higher priority than $i$, i.e., $r_{s}(i)=k$ for each $s\in S$. Let $\succ_{i}$ be any preference ordering such that $i$ has at least $k$ schools acceptable. It is easy to see that $i$’s $k$ most preferred schools constitute a top-priority set and that this is $i$’s smallest top-priority set. By Lemma 11, $i$ has a dominant strategy with $\varphi^{k}$, but not with $\varphi^{k-1}$.
@@ B.4 Efficiency Adjusted Deferred Acceptance mechanism
For any assignment $\mu$, a school $s$ is underdemanded if there is no student $i$ such that $s\succ_{i}\mu(i)$. SEADAM is a multi-round algorithm that works as follows. Let $\Gamma$ be a problem, and let $\Gamma^{1}=\Gamma$. For each round $k\geq 1$, let $\mu^{k}$ be the student-optimal assignment for $\Gamma^{k}$. Let $S^{k}$ be the set of underdemanded schools at $\mu^{k}$. If $S^{k}\neq\emptyset$, for each student $i$, let $\mu^{\ast}(i)=\mu^{k}(i)$ if $\mu^{k}(i)$ is underdemanded. Let $\Gamma^{k+1}$ be the problem $\Gamma^{k}$ reduced by $\mu^{k}(S^{k})$ and proceed to round $k+1$. If $S^{k}=\emptyset$, then $\mu^{\ast}(i)=\mu^{k}$ for each student in the problem $\Gamma^{k}$.
@@
Consider any round $k\geq 1$ and assume that $i$ is not assigned yet at the beginning of round $k$. Clearly, we must have $\varphi_{i}(\Gamma^{k^{\prime}})=s$ for each $k^{\prime}\leq k$. Repeated use of Lemma 15 also implies that $\varphi_{j}(\widehat{\Gamma}^{k^{\prime}})=\varphi_{j}(\Gamma^{k^{\prime}})$ for each $j\neq i$ and each $k^{\prime}\leq k$. If $s$ is an underdemanded school at $\varphi(\Gamma^{k})$, then $i$’s assignment under SEADAM is finalized at round $k$ and we have $\Gamma^{k+1}=\widehat{\Gamma}^{k+1}$. So, at each round $k^{\prime}>k$ of SEADAM the assignments we have $\varphi(\Gamma^{k^{\prime}})=\varphi(\widehat{\Gamma}^{k^{\prime}})$. If $s$ is not an underdemanded school, then $\varphi_{i}(\Gamma^{k+1})=s$ and by Lemma 15, $\varphi_{j}(\widehat{\Gamma}^{k+1})=\varphi_{j}(\Gamma^{k+1})$ for each $j\neq i$. Repeating the argument for steps $k^{\prime}\leq k+1$ yields the desired result.

### 29 Feb 2024 v2

@@ 3.2 Incontestable assignments
Our first question is to characterize incontestable assignments under ex-post canonical information. Clearly, incontestable assignments are necessarily individually rational and non-wasteful. Indeed, given an ex-post canonical information $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q,\mu(i),(|\mu(s)|)_{s\in S}\}$, for any compatible problem $\Gamma=(I,S,\widehat{\succ},r,q)$ and assignment $\widehat{\mu}$ we have ^$\widehat{\succ}_{i}=\succ_{i}$, $\widehat{\mu}(i)=\mu(i)$, and $|\widehat{\mu}(s)|=|\mu(s)|$ for each $s\in S$. So, $s\succ_{i}\mu(i)$ implies $s\succ_{i}\widehat{\mu}(i)$. Also, $s\succ_{i}\mu(i)$ and $|\mu(s)|<q_{s}$ imply $s~{}\widehat{\succ}_{i}~{}\widehat{\mu}(i)$ and $|\widehat{\mu}(s)|<q_{s}$.
@@ 4.1 Incontestable outcomes
Consider the following mechanism $\varphi$. For any profile of preferences $\succ$, $\varphi$ first assigns all students who have a top-priority set (to a school in their top-priority set). Then, $\varphi$ consists of a serial dictatorship with the student $i_{1}$ choosing first if $i_{1}$ does not have a top-priority set. With the preferences we have specified for $i_{1}$, there is no preference profile for students $i_{2}$, $i_{3}$, and $i_{4}$ such that $\varphi$ assigns $i_{1}$ to $s_{3}$. If $i_{2}$ has a top-priority set (which happens whenever she finds at least one school acceptable), $i_{1}$ ends up being assigned to either $s_{1}$ or $s_{2}$. Otherwise, $i_{1}$ is assigned to $s_{1}$.
@@
$\varphi_{i}(\succ,r,q)=s\qquad\Rightarrow\qquad\varphi_{j}(\Gamma)=\varphi_{j}(\Gamma^{\prime})\quad\text{ for each }j\neq i$ (1)
@@
Theorem 2 also establishes an equivalence between the interim perspective and the ex-post perspective (given by the lef-hand side and right-hand side of Eq. (2), respectively). For a student without any top-priority set, any individually rational outcome is a priori (interim) possible, and for any such outcome there exists a preference profile of the other students that rationalizes it (ex-post). Similarly, for a student witha top-priority set, any assignment in that set is a priori (interim) possible, and for any school in that set there exists a compatible problem that rationalizes it (ex-post).
@@ 4.2 Stability and efficiency
Corollary 1 implies that any mechanisms that Pareto improve the Student-Optimal Stable mechanism (e.g., Kesten,’s (2010) efficiency adjusted mechanism) are incontestable. We provide more details on this later in this section.
@@ 4.3 Contestable mechanisms
Because respect for top-priority sets has bite only for the students who have a top-priority set, the set of incontestable assignments is likely to be significantly larger than the set of stable assignments. However, incontestability is not an ‘anything goes’ concept: several well-known assignment mechanisms are contestable. We provide three examples.
@@ 5 Incentives
The main take-away from Theorem 2 is that incontestable and top-top consistent mechanisms are indistinguishable. But that result also turns out to be a tool to analyze such mechanisms. This section aims at illustrating the power of the characterization provided by Theorem 2 to analyze various incentives properties. More precisely, we review in this section three types of standard incentive problems for assignment mechanisms: the existence of safe strategies, maxmin strategies, and manipulability comparisons. Such questions have already been addressed in the literature, but each time for specific mechanisms. We show in this section that thanks to Theorem 2, those properties can be analyzed at once for all incontestable and top-top consistent mechanisms.
@@ 5.1 Safe strategies
Decerf and Van der Linden, (2021) identify a condition under which a student has a ‘safe’ strategy when the mechanism is SOSM. We show here that, thanks to Theorem 2, their result can be easily extended to any incontestable and top-top consistent mechanism. To this end, for a student $i$ and a profile of priority rankings and capacities $(r,q)$, we say that a strategy $\succ_{i}$ is safe under mechanism $\varphi$ if student $i$ is guaranteed to be assigned to a school for any preference profile of the other students. Formally, $\succ_{i}$ is safe for student $i$ if
@@
Proposition 4 has a useful implication, as it permits us to compare how safe a mechanism when varying the constraint imposed on students’ preference lists.
@@ 5.2 Maxmin strategies
Reny, (2022) studies the incentive property of EADAM. [^23] Reny,’s (2022) proposes a different mechanism that EADAM’s, but that turn out to be outcome equivalent to EADAM when all students waive their priorities. That mechanism is no strategyproof, but turns out to still carry some incentives. If students compare two strategies (i.e., preference lists over schools) and compare them with the worst possible outcome (according to the true preferences), then a student can do no better than being truthful. It is worth noting that this belief-free approach, which consists of looking at all possible preference lists of the other students, is shared with our definition of incontestability.
Given a mechanism $\varphi$ and an interim canonical information set $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q\}$, let $w_{i}(\succ_{i}^{\prime}|\succ_{i},\varphi,\mathcal{H}_{i})$ be the worst school according to $\succ_{i}$ when student $i$ submits the preference ordering $\succ_{i}^{\prime}$ when participating to the mechanism. Formally, $w_{i}(\cdot)$ is defined as follows,
@@
Theorem 5 of Reny, (2022) shows that under the priority efficient mechanism (or EADAM) truth-telling is a maxmin optimal strategy for every student. We can easily generalize Reny’s theorem to any incontestable and top-top consistent mechanism.
@@ 5.3 Dominant strategies
Comparisons in terms of dominant strategies inform about the extent to which mechanisms are manipulable. Over the past decades, some school districts reformed their assignment procedures in the objective to reduce their vulnerability to manipulation (Pathak and Sönmez,, 2013). A mechanism that is less vulnerable to manipulations ‘levels the playing field’ because it should a priori penalizes less students who do not strategize well. Understanding to what degree a mechanism is vulnerable to strategic manipulations is of interest for several reasons because strategyproof mechanisms do not necessarily eliminate the incentive to misrepresent one’s preferences when students are constrained, a common practice —see Haeringer and Klijn, (2009). Comparing the manipulability of mechanisms is now a standard exercice. However, this is often done by comparing two explicit mechanisms. We show here that Theorem 2 permits us to compare wholly different mechanisms, for instance, TTC when students are constrained to list at most $k$ schools and SOSM when students are constrained to list at most $k+1$ schools.
@@
Proposition 6.
@@
(i)  The characterization of dominant strategies provided in part (iii) of Proposition 6 does not depend on the specific mechanism. So a student has a dominant strategy in $\varphi^{k}$ if, and only if she has a a dominant strategy in $\psi^{k}$.
(ii)  We first show that $\varphi^{k}$ is less manipulable than $\varphi^{k-1}$. To this end, suppose that student $i$ has a dominant strategy with $\varphi^{k-1}$. By Proposition 6, either $i$ has no more than $k-1$ acceptable schools or $i$’s $k-1$ most preferred schools constitute a top-priority set. These two properties still hold under $\varphi^{k}$, so by Proposition 6 $i$ has a dominant strategy with $\varphi^{k}$.
Consider now the problem $(I,S,\succ,r,q)$, where $q_{s}=1$ for each $s\in S$, and for any two schools $s,s^{\prime}\in S$ and any student $i\in I$, $r_{s}(i)=r_{s^{\prime}}(i)$. Let $i$ be the student for whom exactly $k-1$ other students have higher priority than $i$, i.e., $r_{s}(i)=k$ for each $s\in S$. Let $\succ_{i}$ be any preference ordering such that $i$ has at least $k$ schools acceptable. It is easy to see that $i$’s $k$ most preferred schools constitute a top-priority set and that this is $i$’s smallest top-priority set. By Proposition 6, $i$ has a dominant strategy with $\varphi^{k}$, but not with $\varphi^{k-1}$.
@@ B.4 Efficiency Adjusted Deferred Acceptance mechanism
For any assignment $\mu$, a school $s$ is under-demanded if there is no student $i$ such that $s\succ_{i}\mu(i)$. SEADAM is a multi-round algorithm that works as follows. Let $\Gamma$ be a problem, and let $\Gamma^{1}=\Gamma$. For each round $k\geq 1$, let $\mu^{k}$ be the student-optimal assignment for $\Gamma^{k}$. Let $S^{k}$ be the set of under-demanded schools at $\mu^{k}$. If $S^{k}\neq\emptyset$, for each student $i$, let $\mu^{\ast}(i)=\mu^{k}(i)$ if $\mu^{k}(i)$ is under-demanded. Let $\Gamma^{k+1}$ be the problem $\Gamma^{k}$ reduced by $\mu^{k}(S^{k})$ and proceed to round $k+1$. If $S^{k}=\emptyset$, then $\mu^{\ast}(i)=\mu^{k}$ for each student in the problem $\Gamma^{k}$.
@@
Consider any round $k\geq 1$ and assume that $i$ is not assigned yet at the beginning of round $k$. Clearly, we must have $\varphi_{i}(\Gamma^{k^{\prime}})=s$ for each $k^{\prime}\leq k$. Repeated use of Lemma 3 also implies that $\varphi_{j}(\widehat{\Gamma}^{k^{\prime}})=\varphi_{j}(\Gamma^{k^{\prime}})$ for each $j\neq i$ and each $k^{\prime}\leq k$. If $s$ is an under-demanded school at $\varphi(\Gamma^{k})$, then $i$’s assignment under SEADAM is finalized at round $k$ and we have $\Gamma^{k+1}=\widehat{\Gamma}^{k+1}$. So, at each round $k^{\prime}>k$ of SEADAM the assignments we have $\varphi(\Gamma^{k^{\prime}})=\varphi(\widehat{\Gamma}^{k^{\prime}})$. If $s$ is not an under-demanded school, then $\varphi_{i}(\Gamma^{k+1})=s$ and by Lemma 3, $\varphi_{j}(\widehat{\Gamma}^{k+1})=\varphi_{j}(\Gamma^{k+1})$ for each $j\neq i$. Repeating the argument for steps $k^{\prime}\leq k+1$ yields the desired result.